OpenStudy (anonymous):

Write in standard form of the line that passes through the point (2,-3). its perpendicular. 6x-8y-5=0

OpenStudy (johnweldon1993):

well remember a perpendicular line has the same y intercept....but the negative reciprocal of the slope.... so lets put the equation in slope intercept for a second 6x - 8y - 5 = 0 6x - 8y = 5 -8y = 5 - 6x y = -5/8 + 6/8x written correctly y = 6/8x - 5/8 Okay so thats our line we have.....now we have a point of the other line use the point slope formula for this...knowing that the slope must be the negative resiprocal of this other line so y - (-3) = -8/6(x - 2) y + 3 = -8/6x + 16/6 y = -8/6x + 16/6 - 3 y = -8/6x - .33333333 now check...using the point given....plug into this equation -3 = -8/6(2) -.3333333 -3 = -3 check...so thats your equation for the second line but thats in slope intercept form...and it asks for standard form so y = -8/6x - .33333333 y + 8/6x + .3333333 = 0 p.s .333 could also be written as 1/3 so you have y + 8/6x + 1/3 = 0 Hope I helped :)

OpenStudy (anonymous):

thanks:p

OpenStudy (johnweldon1993):

Anytime :)